# PhishLang A Lightweight, Client-Side Phishing Detection Framework using MobileBERT for Real-Time, Explainable Threat Mitigation

### v2

@@ 7. Client-side extension
Also, recent research (Pourmohamad et al., 2024) has indicated that client-side implementations of in-browser anti-phishing measures, such as Google Safe Browsing, typically rely on simple heuristics and basic machine learning algorithms. These methods often lead to high false positive rates and low detection rates (14%) for threats when not using online blocklists. In contrast, our extension offers an open-source alternative that utilizes a more advanced and flexible model architecture. In Section X of the Appendix, we evaluate the performance and efficiency of our extension across various system configurations. Our findings indicate that the extension uses between 28MB and 247MB of RAM at peak load, which is a small fraction of the memory available in most modern systems. It can also run efficiently on low-end systems with only 1GB of RAM and a single-core processor while maintaining high precision (0.95) and recall (0.96) rates. To enhance performance, the application includes an optional built-in whitelist that users can choose to disable. Users also have the option to report URLs to our PhishLang database, though this feature is completely voluntary and does not prompt for user action. Additionally, users can add detected websites to a personal ’Safe List’ or report URLs as false positives to the PhishLang database if necessary. Figure 5 provides an illustration of the client-side extension in operation.
The Explainable Blocklisting feature is also available, though it is disabled by default. To enable this feature, users must provide their OpenAI key to access the explainable functionality, as detailed in Section 6. While this requires sharing the parsed representation of the website being visited, OpenAI allows users to prevent their provided prompts or generated responses used from being used for training or other purposes (OpenAI, 2024). We plan to conduct future research to offer this explainable feature on the client side as well, especially as more resource-efficient open-domain chatbots like Gemma become available (Google, 2024). However, these chatbots are not yet lightweight enough for practical local use on low to medium-end systems.
@@ 8. Conclusion
In this work, we introduced PhishLang, a novel phishing detection framework leveraging the MobileBERT architecture to identify phishing websites by comparing source codes with established phishing patterns. PhishLang outperforms various machine learning and deep learning models, reducing detection time and resources while showing strong resilience against adversarial attacks. In real-time, it identified several new phishing websites, including evasive attacks that were missed by popular anti-phishing tools. Reporting these threats improved detection rates, demonstrating PhishLang’s impact. We integrated PhishLang with GPT-3.5T for ”Explainable Blocklisting,” providing contextual explanations for phishing alerts, enhancing user understanding. We also implemented PhishLang as the first open-source client-side browser extension that does not require an external blocklist.

### v3

@@ 6. Client-side extension
Also, recent research (Pourmohamad et al., 2024) has indicated that client-side implementations of in-browser anti-phishing measures, such as Google Safe Browsing, typically rely on simple heuristics and basic machine learning algorithms. These methods often lead to high false positive rates and low detection rates (14%) for threats when not using online blocklists. In contrast, our extension offers an open-source alternative that utilizes a more advanced and flexible model architecture. In Section 6.1, we evaluate the performance and efficiency of our extension across various system configurations. Our findings indicate that the extension uses between 28MB and 247MB of RAM at peak load, which is a small fraction of the memory available in most modern systems. It can also run efficiently on low-end systems with only 1GB of RAM and a single-core processor while maintaining high precision (0.95) and recall (0.96) rates. To enhance performance, the application includes an optional built-in whitelist that users can choose to disable. Users also have the option to report URLs to our PhishLang database, though this feature is completely voluntary and does not prompt for user action. Additionally, users can add detected websites to a personal ’Safe List’ or report URLs as false positives to the PhishLang database if necessary. Figure 8 provides an illustration of the client-side extension in operation.
@@
6.1. Evaluating the performance of our client-side extension
We evaluate 2,000 verified phishing websites from PhishTank, comparing the performance of PhishLang on four distinct system configurations. The first configuration is our experimental setup, featuring an Intel Xeon W Processor with 184GB of RAM and 4x NVIDIA A5000 GPUs, running Ubuntu 22.04 LTS, with the model running natively. The second configuration is a mid-tier setup, consisting of an Intel Core i5 13th Gen processor, 8GB of RAM, no GPU, and running Windows 11 Build 22631. The third configuration is a low-tier setup, using an Intel Celeron N4500 processor, 2GB of RAM, no GPU, and running Ubuntu 24.04 LTS. The final configuration involves a virtual machine using VMware Player emulated with 2 core, 4GB of RAM, no GPU acceleration, and running Windows 11 Build 22631. These various configurations were chosen to assess PhishLang’s performance across a spectrum of hardware configurations that reflect typical end-user environments.
Table 9 presents the performance of PhishLang in terms of efficiency, memory usage, and inference time. Our findings show that the client-side application running on both Ubuntu and Windows 11 achieves the same efficiency as the model running on our experimental server, indicating no performance loss when PhishLang operates on the client-side. Additionally, all configurations exhibit similar memory usage. Memory usage was calculated by considering the system service (local server) and the memory consumed by the browser extension. It is noteworthy that the experimental setup has lower RAM usage since it runs the base model without any browser extension overhead. Regarding inference time, the low-tier and virtual machine configurations, despite having weaker processors, demonstrate only a slight decrease in speed compared to the experimental and mid-tier setups. All configurations maintain a median inference time of under one second. This suggests that our application, powered by the lightweight MobileBERT architecture, offers rapid inference regardless of the system configuration. It is important to note that the reported inference time excludes the time required to fetch the website, as server congestion or intentional delays by attackers can significantly affect website loading times.
Configuration Precision/Recall Memory usage Inference time (in secs)
Min Max Median Min Max Median
Experimental 0.95/0.96 28MB 215MB 76MB 0.08 1.09 0.44
Medium-end 0.95/0.96 51MB 308MB 93MB 0.13 1.47 0.59
Low-end 0.95/0.96 54MB 247MB 89MB 0.18 1.95 0.73
Virtual Machine 0.95/0.96 68MB 231MB 102MB 0.11 2.43 0.81
Table 9. Performance of the PhishLang app for various system configurations
Trade-offs:
Currently, running PhishLang as a client-side application requires downloading TensorFlow and PyTorch, which the installer automatically handles if they are not already installed. This setup can take up to 2.5GB of storage space on the user’s machine. However, since storage has become significantly cheaper over the years, it is reasonable to assume that most end-user systems can accommodate these requirements. An alternative to reduce storage space is to implement the model using TensorFlowJS directly within the browser extension. However, this approach has significant drawbacks. Running the model in the browser with TensorFlowJS may lead to slower performance compared to running it natively as a local server application, as browser-based implementations (especially on lower-end configurations) do not leverage hardware acceleration as effectively. Additionally, browser environments can be inconsistent across various devices and operating systems, which might lead to compatibility and stability issues. This inconsistency could result in the extension not functioning uniformly across all browsers, potentially affecting the reliability of the phishing detection system. Moreover, implementing the extension separately for different browsers would add to the development complexity. In contrast, a local server-based implementation would ensure broader compatibility, as it only requires basic communication capabilities between the extension and the local server, a feature supported by all major browsers.
@@ 7. Conclusion
In this work, we introduced PhishLang, a novel phishing detection framework leveraging the MobileBERT architecture to identify phishing websites by comparing source codes with established phishing patterns. PhishLang outperforms various machine learning and deep learning models, reducing detection time and resources while showing strong resilience against adversarial attacks. In real-time, it identified several new phishing websites, including evasive attacks that were missed by popular anti-phishing tools. Reporting these threats improved detection rates, demonstrating PhishLang’s impact. We also implemented PhishLang as the first open-source client-side browser extension that does not require an external blocklist.